# Parametric Phase Modulation in Superconducting Circuits

### v1

@@ Appendix B Spectroscopy of parametrically modulated qubits
Figure 8: Dependence of first-order ($n=1$) sideband coupling on the parametric drive amplitude $\tilde{\Phi}$. The data are for scenarios where either $Q_{1}$ or $Q_{2}$ is independently driven, while the other qubit is maintained at a static DC flux bias. (a) Effective coupling strength $2g_{\mathrm{eff}}^{1}/2\pi$ versus parametric amplitude $\tilde{\Phi}$, extracted from the period of population oscillations between the $|10\rangle$ and $|01\rangle$ states. (b) Parametric frequency $\omega_{p}/2\pi$ required to maintain the sideband resonance condition, plotted against $\tilde{\Phi}$. (c) The resulting resonant frequency deviation, $|\Delta_{p}|=|\omega_{p}(\tilde{\Phi})-\omega_{p}(0)|$, plotted against the achieved effective coupling strength $2g_{\mathrm{eff}}^{1}/2\pi$. This panel highlights the significant frequency adjustment required when tuning the coupling via amplitude.
Appendix C Flux pulse transfer function
@@
Appendix D Coupling strength of different parametric amplitudes
@@
To demonstrate the parametric coupling strengths as parametric amplitudes increase, we focus on the first-order sideband resonance of two qubits at off-sweet spots. In the linear region, where the qubit frequency versus external flux is nearly linear, the relation between parametric coupling strengths and amplitudes follows the first-order Bessel function of the first kind [15]. However, due to the nonlinearity of transmon qubits, the first-order sideband resonant frequency deviates as the parametric amplitude increases. Additionally, the qubit frequency excursions during modulation, $\epsilon_{p}$, become increasingly unpredictable (see Appendix E for details).
@@
We bias two qubits at off-sweet spots, with $\bar{\Phi}_{1}=0.119\Phi_{0}$ and $\bar{\Phi}_{2}=0.1025\Phi_{0}$, to measure the effective coupling strength as a function of the parametric amplitude $\tilde{\Phi}$. As the amplitude increases, the optimal parametric flux frequency for achieving first-order sideband resonance can shift, which can be predicted using a three-tone spectroscopy experiment. In the experiment, we first prepare the $|10\rangle$ state and then impose a parametric flux pulse on $Q_{1}$ ($Q_{2}$) while keeping $Q_{2}$ ($Q_{1}$) at a static DC flux bias. We measure the population oscillations of the $|10\rangle$ state, and from the oscillations versus parametric frequencies, we can extract both the corresponding effective coupling strengths $2g_{\mathrm{eff}}^{1}$ (shown in Fig. 8(a)) and the parametric resonance frequencies $\omega_{p}$ (shown in Fig. 8(b)). This amplitude-based tuning method reveals a significant practical challenge: the nonlinearity of the qubit’s frequency response to external flux causes the required resonant drive frequency, $\omega_{p}$, to shift substantially as the parametric amplitude $\tilde{\Phi}$ is varied. As illustrated in Fig. 8(c), this frequency deviation, defined as $|\Delta_{p}|=|\omega_{p}(\tilde{\Phi})-\omega_{p}(0)|$, can exceed $100$ MHz to achieve a desired coupling strength. Such a large co-variation necessitates cumbersome, multi-parameter calibration routines. In stark contrast, our phase modulation scheme provides a decoupled control knob, allowing the coupling strength to be tuned without inducing these parasitic frequency shifts, thereby greatly simplifying the calibration process.
@@
By combining the transfer function and pre-calibrated three-tone spectroscopy, we can effectively explore parametric coupling under large frequency deviations, where the flux-to-frequency transduction is highly nonlinear.
Figure 9: Comparison of Taylor expansion and Fourier series for calculating components of the parametrically modulated qubit frequency. (a) DC frequency shift of the qubit under parametric modulation, shown as a function of the expansion order $n$. Results are presented for modulation at the sweet spot ($\bar{\Phi}=0,\tilde{\Phi}=0.4\Phi_{0}$; teal lines) and an off-sweet spot ($\bar{\Phi}=0.15\Phi_{0},\tilde{\Phi}=0.3\Phi_{0}$; fuchsia lines). Solid lines represent calculations from Taylor expansion, while dashed lines show the corresponding Fourier series components. (b) Amplitude of the dominant AC harmonic component of the modulated qubit frequency versus expansion order $n$. For the sweet spot (teal lines), this is the second harmonic (frequency $2\omega_{p}$ component); for the off-sweet spot (fuchsia lines), it is the first harmonic (frequency $\omega_{p}$ component). Taylor expansion results (solid lines) are shown approaching the Fourier series values (dashed lines) as $n$ increases. All calculations use transmon qubit parameters $E_{C}/2\pi=240$ MHz and $E_{J1}/2\pi=E_{J2}/2\pi=8.286$ GHz.
Appendix E Numerical comparison between the Taylor expansion and the Fourier series
Figure 10: Average deviations of Taylor expansion and Fourier series approximations from an analytical solution for the parametrically modulated qubit frequency $\omega(t)$. (a) Average deviation over one modulation period from the analytical solution for calculations at the sweet spot, plotted as a function of increasing expansion order $n$. (b) Corresponding average deviation for calculations at the off-sweet spot. In both (a) and (b), results for the Fourier series are shown as solid lines with circle markers, while results for the Taylor expansion are shown as solid lines with triangle-up markers. Insets: Time evolution of the modulated qubit frequency, $\omega(t)$, over one period, comparing the analytical solution (black solid line) with approximations from Fourier series (fuchsia dashed line) and Taylor expansion (teal dotted line) at expansion orders $n=3$ and $n=10$. All physical and modulation parameters are identical to those used in Fig. 9.
Using the power-reduction formulae, we can rewrite the Taylor expansion (12) as
$\begin{split}\omega(t)&=\left[\omega(\bar{\Phi})+\sum_{n=1}^{\infty}\frac{\tilde{\Phi}^{n}}{2^{n}n!}\left.\frac{\partial^{n}\omega}{\partial\Phi^{n}}\right|_{\bar{\Phi}}\delta_{{n\bmod 2},0}\binom{n}{\lfloor\frac{n}{2}\rfloor}\right]\\ &+\sum_{n=1}^{\infty}\frac{2\tilde{\Phi}^{n}}{2^{n}n!}\left.\frac{\partial^{n}\omega}{\partial\Phi^{n}}\right|_{\bar{\Phi}}\sum_{k=0}^{\lfloor\frac{n-1}{2}\rfloor}\binom{n}{k}\cos((n-2k)\omega_{p}t).\end{split}$ (13)
As the expansion order $n$ increases, more harmonics appear, improving accuracy. We choose typical parameters $\bar{\Phi}=0,\tilde{\Phi}=0.4\Phi_{0}$ at sweet spots and $\bar{\Phi}=0.15\Phi_{0},\tilde{\Phi}=0.3\Phi_{0}$ at off-sweet spots as examples.
As shown in Fig. 9 (a), the DC shifts depend on the even-order terms of the Taylor expansion. The main shift can be approximated by the second-order ($n=2$) term. As the Taylor order $n$ increases, the DC component approaches the zeroth-order of the Fourier series, corresponding to the time-averaged frequency for one period. The DC component results from the nonlinearity in transmon qubits and can also be derived using the charge and flux operators, as shown in [42], rather than the analytical equation provided above.
Figure 11: Average deviations of Taylor expansion and Fourier series approximations (at a fixed expansion order of $n=3$) from an analytical solution for the parametrically modulated qubit frequency, shown as the parametric amplitude $\tilde{\Phi}$ increases. Solid and dashed lines represent calculations for the qubit at its sweet and off-sweet spots, respectively. Fuchsia lines correspond to the Fourier series approximation, and teal lines correspond to the Taylor expansion approximation. All other physical and modulation parameters are identical to those specified in Fig. 9.
As shown in Fig. 9 (b), we respectively estimate the second harmonic at the sweet spot and the first harmonic at the off-sweet spot, which are used to theoretically determine the parametric coupling strength at weak parametric modulation amplitudes. The main harmonic components are obtained by considering the Taylor expansion at order $n=2$ (the sweet spot) and at order $n=3$ (the off-sweet spot). At sweet spots, the qubit frequency oscillates only at even harmonics of the parametric frequency since all odd-order derivatives vanish. No significant difference is observed for Taylor expansion orders $n>6$ at sweet and off-sweet spots.
The time-dependent frequency of qubits under parametric modulation can be approximated using both the Taylor expansion and the Fourier series. The average deviations over a period between these methods and the analytical equation decrease as the expansion order $n$ increases at sweet and off-sweet spots, as shown in Fig. 10. The insert figures show the time evolution of qubit frequencies over one period at orders $n=3$ and $n=10$. At $n=3$, the Taylor expansion slightly deviates, while no visible deviations occur at higher orders. At sweet spots, the Fourier spectrum exhibits a stepped shape due to the vanishing of odd Fourier coefficients [9]. Average deviations are lower at sweet spots compared to off-sweet spots, due to the symmetry of flux-to-frequency transduction. Both methods provide good approximations of qubit time-dependent frequencies and coupling strength calculations for weak parametric amplitudes, but the accuracy decreases as the parametric amplitude increases, as shown in Fig. 11. At sweet spots, both methods yield lower deviations due to the symmetry of flux-to-frequency transduction, which eliminates odd harmonics.
@@ Appendix E Numerical comparison between the Taylor expansion and the Fourier series
Appendix F Parametric-resonance phase modulation
Parametric resonance is a technique for realizing fast two-qubit entangling gates and mitigating issues such as frequency collisions in superconducting qubit systems [19]. The phase modulation approach detailed in the main text can also be effectively applied to these parametric resonance interactions (i.e., the zeroth-order ($n=0$) sideband coupling), and this application is demonstrated in Fig. 12.
The experimental procedure for achieving phase-modulated coupling via parametric resonance is analogous to that described for the first-order sideband coupling in the main text. The primary distinction lies in the activation condition: for zeroth-order sideband coupling (parametric resonance), the interaction is engaged by tuning the two parametrically modulated qubits such that their time-averaged frequencies become resonant (i.e., $\bar{\omega}_{1}=\bar{\omega}_{2}$). This contrasts with first-order sideband couplings, where resonance is achieved when the effective detuning matches a non-zero integer multiple of the parametric flux frequency (e.g., $\Delta+\omega_{p}=0$).
@@ Appendix F Parametric-resonance phase modulation
Appendix G Coherence under parametric phase modulation
Coherence is central to high-fidelity operations. For qubits under parametric modulation, coherence is influenced by both their intrinsic properties and the surrounding environment. The flux-noise spectra of qubits determine their dephasing rates, and while modulation can sometimes increase dephasing due to effects like multiplicative 1/f noise. Beyond flux noise, the qubit’s environment, particularly two-level system (TLS) defects, is a common source of decoherence. Theoretical work has shown that frequency modulation can help stabilize relaxation rates and mitigate specific dephasing mechanisms arising from TLSs [31, 32]. The stability and quality of qubits can also be affected by slow periodic frequency modulation [33].
For the flux noise, there are powerful mitigation strategies. For instance, operating at specific modulation parameters can create “dynamical sweet spots” that are insensitive to 1/f flux noise [24, 14], a technique that has been successfully used to demonstrate high-fidelity CZ gates [26]. The dephasing rate $\Gamma_{\phi,1/f}$ under modulation can be described as
$\Gamma_{\phi,1/f}=\lambda\left|\frac{\partial\bar{\omega}}{\partial\tilde{\Phi}}\right|A_{ac,1/f},$ (14)
where $\lambda$ and $A_{ac,1/f}$ represent the noise parameter and amplitude due to 1/f noise [24]. This concept can be extended; for example, two-tone parametric modulation can create a continuum of dynamical sweet spots, expanding the range of flux-noise-robust operating frequencies [25, 29]. The benefits of parametric drives have also been explored for enhancing coherence in bichromatically driven Floquet qubits [30]. Parametric modulation can also be applied to fluxonium qubits to protect qubits from 1/f noise and enhance coherence times [28, 27]. Indeed, parametric modulation can be viewed as a continuous version of the dynamical decoupling scheme to realize high-fidelity Controlled-Z Gates which can greatly reduce the qubit dephasing [23].
Coherence degradation is a known challenge for some hardware-based tunable coupler schemes [57, 58]. Such degradation often occurs when the system is tuned into a highly hybridized regime, where the coherence of the qubits can be limited by that of the coupler which can be explained using a two-spin Hamiltonian [59] or the resulting dressed states become more sensitive to flux noise. The effective decoherence rates in such systems can be modeled by the participation ratios of the uncoupled modes [60].
In our experiment, we have designed our scheme to specifically enhance coherence. We choose two qubits with a coupler in our chip and other irrelevant couplers are biased to idle points with zero couplings. The chosen hardware coupler is biased to a fixed operating point that maintains a sufficient effective coupling strength while ensuring the qubits are only weakly hybridized with it. Crucially, our working points for the qubits are at or near their DC sweet spots, leveraging their inherent insensitivity to flux noise [49]. Consequently, the dual parametric drives, which are simultaneously applied to the two adjacent qubits, are designed to inherit the robustness of parametric modulation against pulse distortions and noise. The introduction of a second, synchronous drive primarily adds a new degree of control—the relative phase—without fundamentally altering the underlying noise-protection mechanisms.
@@ Appendix H Stability and scalability of parametric phase modulation
In our manuscript, we have defined the main variable for the relative phase between the two drives as $\delta\phi_{p}$. Fluctuations in this quantity will translate into fluctuations in the coupling strength, $\delta g_{\text{phase}}^{n}$. To avoid notational ambiguity, we will denote a small, random fluctuation (i.e., jitter) in our relative phase variable as $j_{\phi}$.
@@
Fluctuations of absolute parametric phases usually arise from the electronic devices, lines, and the thermal and electromagnetic environment. The dominant source of these fluctuations is typically the relative phase jitter between the output channels of the arbitrary waveform generator (AWG). For modern high-performance AWGs used in quantum control, a typical timing jitter of $\sim 5$ for a $100$ MHz parametric drive results in a phase fluctuation of $j_{\phi}\sim 0.003$ rad [61].
@@
We have performed a calculation based on our theoretical model to estimate the impact of this jitter. The fluctuation in coupling strength can be approximated as $\delta g_{\text{phase }}^{n}\approx\left|\frac{dg_{\text{phase }}^{n}}{d\left(\delta\phi_{p}\right)}\right|j_{\phi}$. Using the chain rule and the properties of Bessel functions, the sensitivity is given by:
@@
$\begin{split}\left|\frac{dg_{\text{phase }}^{n}}{d\left(\delta\phi_{p}\right)}\right|&=\left|\frac{d\left(gJ_{n}(A)\right)}{dA}\frac{dA}{d\left(\delta\phi_{p}\right)}\right|\\ &=\left|\frac{g}{2}\left[J_{n-1}(A)-J_{n+1}(A)\right]\frac{dA}{d\left(\delta\phi_{p}\right)}\right|\\ &=\left|\frac{g}{2}\left[J_{n-1}(A)-J_{n+1}(A)\right]\frac{\epsilon_{p1}\epsilon_{p2}}{A\omega_{p}^{2}}\sin\left(\delta\phi_{p}\right)\right|\end{split}$ (15)
@@
where $A$ is the phase-dependent argument of the Bessel function. We evaluated this expression at the most sensitive point of our experimental data in Fig. 2 (a) yielding a sensitivity of $\left|\frac{dg_{\text{phase }}^{n}}{d\left(\delta\phi_{p}\right)}\right|\approx 22$ MHz/rad. The resulting fluctuation in coupling strength is therefore $\delta g_{\text{phase }}^{n}\approx 0.066$ MHz. Given that our demonstrated coupling strengths are on the order of several MHz, this corresponds to a relative fluctuation of less than $1\%$ at the most sensitive point. This level of fluctuation is negligible for our experiments and does not require frequent recalibration.
@@
We utilize a two-qubit system with a tunable coupler to demonstrate our scheme and these qubits are also tunable which is common in large-scale quantum simulation and computation. Regarding wiring complexity, our scheme is highly scalable as it requires no additional physical hardware elements. The dual parametric drives are delivered through the standard on-chip flux bias lines that are already present for individual qubit control in state-of-the-art multi-qubit processors. Therefore, it introduces no new hardware overhead.
@@
Regarding crosstalk, we agree that flux crosstalk is a key challenge in scaling superconducting processors. In our scheme, since the parametric drives are applied locally via these individual flux lines, the crosstalk considerations are fundamentally the same as those in any multi-qubit system employing simultaneous flux control. We therefore expect that established DC and AC flux crosstalk mitigation techniques, such as those described in Ref. [18], can be directly applied to characterize and suppress these effects.
@@
In summary, we believe our technique is well-suited for practical applications in larger systems because it adds a powerful layer of control without introducing new hardware or wiring complexity and is compatible with existing solutions for crosstalk management. We also note that it may even offer new strategies for its mitigation.

### v2

@@ Appendix E Flux pulse transfer function
Appendix F Coupling strength of different parametric amplitudes
Figure 11: Dependence of first-order ($n=1$) sideband coupling on the parametric drive amplitude $\tilde{\Phi}$. The data are for scenarios where either $Q_{1}$ or $Q_{2}$ is independently driven, while the other qubit is maintained at a static dc flux bias. (a) Effective coupling strength $2g_{\mathrm{eff}}^{1}/2\pi$ versus parametric amplitude $\tilde{\Phi}$, extracted from the period of population oscillations between the $|10\rangle$ and $|01\rangle$ states. (b) Parametric frequency $\omega_{p}/2\pi$ required to maintain the sideband resonance condition, plotted against $\tilde{\Phi}$. (c) The resulting resonant frequency deviation, $|\Delta_{p}|=|\omega_{p}(\tilde{\Phi})-\omega_{p}(0)|$, plotted against the achieved effective coupling strength $2g_{\mathrm{eff}}^{1}/2\pi$. Panel (c) highlights the significant frequency adjustment required when tuning the coupling via amplitude.
@@ Appendix F Coupling strength of different parametric amplitudes
To demonstrate the parametric coupling strengths as parametric amplitudes increase, we focus on the first-order sideband resonance of two qubits at off-sweet spots. In the linear region, where the qubit frequency versus external flux is nearly linear, the relation between parametric coupling strengths and amplitudes follows the first-order Bessel function of the first kind [15]. However, due to the nonlinearity of transmon qubits, the first-order sideband resonant frequency deviates as the parametric amplitude increases. Additionally, the qubit frequency excursions during modulation, $\epsilon_{p}$, become increasingly unpredictable (see Appendix C for details).
@@
We bias two qubits at off-sweet spots, with $\bar{\Phi}_{1}=0.119\Phi_{0}$ and $\bar{\Phi}_{2}=0.1025\Phi_{0}$, to measure the effective coupling strength as a function of the parametric amplitude $\tilde{\Phi}$. As the amplitude increases, the optimal parametric flux frequency for achieving first-order sideband resonance can shift, which can be predicted using a three-tone spectroscopy experiment. In the experiment, we first prepare the $|10\rangle$ state and then impose a parametric flux pulse on $Q_{1}$ ($Q_{2}$) while keeping $Q_{2}$ ($Q_{1}$) at a static dc flux bias. We measure the population oscillations of the $|10\rangle$ state, and from the oscillations versus parametric frequencies, we can extract both the corresponding effective coupling strengths $2g_{\mathrm{eff}}^{1}$ [shown in Fig. 11(a)] and the parametric-resonance frequencies $\omega_{p}$ [shown in Fig. 11(b)]. This amplitude-based tuning method reveals a significant practical challenge: the nonlinearity of the qubit’s frequency response to external flux causes the required resonant drive frequency, $\omega_{p}$, to shift substantially as the parametric amplitude $\tilde{\Phi}$ is varied. As illustrated in Fig. 11(c), this frequency deviation, defined as $|\Delta_{p}|=|\omega_{p}(\tilde{\Phi})-\omega_{p}(0)|$, can exceed $100$ MHz to achieve a desired coupling strength. Such a large covariation necessitates cumbersome, multiparameter calibration routines. In stark contrast, our phase-modulation scheme provides a decoupled control knob, allowing the coupling strength to be tuned without inducing these parasitic frequency shifts, thereby greatly simplifying the calibration process.
@@
By combining the transfer function and precalibrated three-tone spectroscopy, we can effectively explore parametric coupling under large frequency deviations, where the flux-to-frequency transduction is highly nonlinear.
Appendix G Parametric-resonance phase modulation
@@
Parametric resonance is a technique for realizing fast two-qubit entangling gates and mitigating issues such as frequency collisions in superconducting qubit systems [19]. The phase-modulation approach detailed in the main text can also be effectively applied to these parametric-resonance interactions (i.e., the zeroth-order ($n=0$) sideband coupling), and this application is demonstrated in Fig. 12.
The experimental procedure for achieving phase-modulated coupling via parametric resonance is analogous to that described for the first-order sideband coupling in the main text. The primary distinction lies in the activation condition: for zeroth-order sideband coupling (parametric resonance), the interaction is engaged by tuning the two parametrically modulated qubits such that their time-averaged frequencies become resonant (i.e., $\bar{\omega}_{1}=\bar{\omega}_{2}$). This contrasts with first-order sideband couplings, where resonance is achieved when the effective detuning matches a nonzero integer multiple of the parametric flux frequency (e.g., $\Delta+\omega_{p}=0$).
@@ Appendix H Stability and scalability of parametric phase modulation
In this paper, we have defined the main variable for the relative phase between the two drives as $\delta\phi_{p}$. Fluctuations in this quantity will translate into fluctuations in the coupling strength, $\delta g_{\text{phase}}^{n}$. To avoid notational ambiguity, we will denote a small, random fluctuation (i.e., jitter) in our relative phase variable as $j_{\phi}$.
@@
Fluctuations of absolute parametric phases usually arise from electronic devices, lines, and the thermal and electromagnetic environment. The dominant source of these fluctuations is typically the relative phase jitter between the output channels of the arbitrary waveform generator. For the modern high-performance AWGs used in quantum control, a typical timing jitter of $\sim 5$ ns for a 100100- MHz parametric drive results in a phase fluctuation of $j_{\phi}\sim 0.003$ rad [62].
@@
We have performed a calculation based on our theoretical model to estimate the impact of this jitter. The fluctuation in coupling strength can be approximated as $\delta g_{\text{phase }}^{n}\approx|{dg_{\text{phase }}^{n}}/{d\left(\delta\phi_{p}\right)}|j_{\phi}$. Using the chain rule and the properties of Bessel functions, the sensitivity is given by
@@
$\begin{split}\left|\frac{dg_{\text{phase }}^{n}}{d\left(\delta\phi_{p}\right)}\right|&=\left|\frac{d\left(gJ_{n}(A)\right)}{dA}\frac{dA}{d\left(\delta\phi_{p}\right)}\right|\\ &=\left|\frac{g}{2}\left[J_{n-1}(A)-J_{n+1}(A)\right]\frac{dA}{d\left(\delta\phi_{p}\right)}\right|\\ &=\left|\frac{g}{2}\left[J_{n-1}(A)-J_{n+1}(A)\right]\frac{\epsilon_{p1}\epsilon_{p2}}{A\omega_{p}^{2}}\sin\left(\delta\phi_{p}\right)\right|,\end{split}$ (15)
@@
where $A$ is the phase-dependent argument of the Bessel function. We evaluated this expression at the most sensitive point of our experimental data in Fig. 2 (a) yielding a sensitivity of $|{dg_{\text{phase }}^{n}}/{d\left(\delta\phi_{p}\right)}|\approx 22$ MHz/rad. The resulting fluctuation in coupling strength is therefore $\delta g_{\text{phase }}^{n}\approx 0.066$ MHz. Given that our demonstrated coupling strengths are on the order of several megahertz, this corresponds to a relative fluctuation of less than $1\%$ at the most sensitive point. This level of fluctuation is negligible for our experiments and does not require frequent recalibration.
@@
We utilize a two-qubit system with a tunable coupler to demonstrate our scheme and these qubits are also tunable which is common in large-scale quantum simulation and computation. Regarding wiring complexity, our scheme is highly scalable as it requires no additional physical hardware elements. The dual parametric drives are delivered through the standard on-chip flux bias lines that are already present for individual qubit control in state-of-the-art multiqubit processors. Therefore, it introduces no new hardware overhead.
@@
Regarding crosstalk, we agree that flux crosstalk is a key challenge in scaling superconducting processors. In our scheme, since the parametric drives are applied locally via these individual flux lines, the crosstalk considerations are fundamentally the same as those in any multiqubit system employing simultaneous flux control. We therefore expect that established dc and ac flux crosstalk mitigation techniques, such as those described in Ref. [18], can be directly applied to characterize and suppress these effects.
@@
In summary, we believe our technique is well suited for practical applications in larger systems because it adds a powerful layer of control without introducing new hardware or wiring complexity and is compatible with existing solutions for crosstalk management. We also note that it may even offer new strategies for its mitigation.